OpenStudy (anonymous):

critical points of x-15x^(1/3)

OpenStudy (anonymous):

set \[1-\frac{5}{\sqrt[3]{x^2}}=0\] and solve

OpenStudy (anonymous):

also since the derivative is undefined if \(x=0\) you have \(0\) as a critical point as well

OpenStudy (anonymous):

i'm not getting it right

OpenStudy (anonymous):

are you stuck at solving for \(x\)?

OpenStudy (anonymous):

or at finding the derivative?

OpenStudy (anonymous):

solving for x

OpenStudy (anonymous):

then you are thinking way too hard take two steps or three

OpenStudy (anonymous):

\[1-\frac{5}{\sqrt[3]{x^2}}=0\\ \frac{5}{\sqrt[3]{x^2}}=1\\ \sqrt[3]{x^2}=5\] is a start

OpenStudy (anonymous):

yeah, I get that far

OpenStudy (anonymous):

oh then if \[\sqrt[3]{x^2}=5\] you have \[x=\sqrt{5^3}\]

OpenStudy (anonymous):

yeah, that's what I thought but its saying its wrong

OpenStudy (anonymous):

it is right

OpenStudy (anonymous):

also 0

OpenStudy (anonymous):

I put both and says that it is wrong. That's why I am so lost about it

OpenStudy (anonymous):

hmm i think it is right maybe it wants a decimal? maybe it wants a decimal to more places? can you post a screenshot?

OpenStudy (anonymous):

hmm i don't see the instructions, but i am lost pretty positive the answer is right

OpenStudy (anonymous):

unless for some reason they want you to write \[5\sqrt5\] but that is awful picky

OpenStudy (anonymous):

that isn't right either

OpenStudy (anonymous):

I figured it out. it was 0, 5sqrt5 and -5sqrt5